the sum of three numbers is 60. one number is five more than a second number. it is also twice the third. find the numbers
Answer by macston(5194)

.
F=first number=; S=second number=(1/5)F; T=third number=(1/2)F
.
F+S+T=60 Substitute for S and T.
F+(1/5)F+(1/2)F=60 Find common denominator and add.
(10/10)F+(2/10)F+(5/10)F=60
(17/10)F=60 Multiply each side by (10/17)
F=600/17=35.29 ANSWER 1: The first number is 35.29.
S=(1/5)F=7.06 ANSWER 2: The second number is 7.06.
T=(1/2)F=17.65 ANSWER 3: The third number is 17.65.
.
CHECK:
F+S+T=60
35.29+7.06+17.65=60
60=60